Indicate the equation of the given line in standard form.
The line through point (-3, 4) and perpendicular to a line that has slope 2/5

Found 2 solutions by sofiyac, amalm06:
Answer by sofiyac(983)
Answer by amalm06(224)

If two non-vertical lines are perpendicular, then the product of their slopes is -1.
Therefore, the slope of the line in question is -5/2.
And the equation of the line in question can be written:
4=(-3)(-5/2)+b so that b=-7/2
Therefore, the equation of the line in standard form is:
y+(5/2)x=-7/2 (Answer)